A 6-month forward contract for corn exists with a price of $1.70 per bushel. If Farmer Jayne decides to hedge her 20,000 bushels of corn with the forward contract, what is her profit or loss if spot prices are $1.65 or $1.80 when she sells her crop in 6 months? Her total costs are $33,000. 5. S0 loss or $3,000 loss $1,000 gain or $1,000 gain A) $1,000 gain or $1,000 loss C) D) B) S0 gain or $3,000 gain Two 6-month corn put options are available. The strike prices are $1.80 and $1.75 with premiums of S0.14 and $0.12 (per bushel), respectively. Average total costs are $1.65 per bushel to produce and 6-month interest rates are 4.0%. Farmer Jayne wishes to hedge 20,000 bushels output for 6 months. What is the highest profit or minimum loss that can be achieved using either one of the two options if the spot price in 6 months is $1.70 per bushel? 6. A) $88 loss B) $88 gain c) $496 loss D) $496 gain Corn call options with a S1.70 strike price are trading for a So.15 premium. Farmer Jayne decides to hedge her 20,000 bushels of corn by selling short call options. Six-month interest rates are 4.0% and she plans to close her position and sell her corn in 6 months. Her average total cost for corn is $1.65 per bushel. What is her profit or loss if spot prices are $1.60 per bushel when she closes her position? 7. A) B) $1,000 loss $2,000 gain C) $2,120 loss D) $2,120 gain
Answer #1

Answer a)

Gain = forward Price * 20000 - cost = $1.7*20000 -$33000 = $1000

Loss = 0

Option D : $1000 gain or $1000 gain.

Answer 2) Put option provides right to sell,

Required price = cost + interest = 1.65 * (1+0.04) = 1.72

In the case , profit or loss = Strike price - required Price - Cost of Put

For 1st Put , profit or loss = ($1.80 - $1.72 - $0.14 ) * 20,000 = -$1200

For 2nd Put, profit or loss = ($1.75 - $1.72- $0.12 ) * 20,000 = -$1800

Althought None in the option :

Answer 3)  Short Call ,obliged you to buy the underlying asset at a fixed price in the future.

Call profit = premium recieved = O.15*20000 =3000

Change in price = 1.70- 1.60 *20,000 = 2000.

Loss = 3000-2000 = 1000 loss ,

Option A.
Answer #2

Q - 5

His net profit of loss = Gail / Loss on sale + Gain / Loss on the future position = (Spot price x Quantity - Cost) + (Forward rate - spot rate) x quantity

Cost = $ 33,000; Quantity = 20,000 bushels; Forward rate = $ 1.70 / bushel

When spot rate = $ 1.65 / bushel then

Her profit or loss = (Spot price x Quantity - Cost) + (Forward rate - spot rate) x quantity = (1.65 x 20,000 - 33,000) + (1.70 - 1.65) x 20,000 = $ 1,000 gain

When spot rate = $ 1.80 / bushel then

Her profit or loss = (Spot price x Quantity - Cost) + (Forward rate - spot rate) x quantity = (1.80 x 20,000 - 33,000) + (1.70 - 1.80) x 20,000 = $ 1,000 gain

Hence, the correct answer is option D. $1,000 gain or $1,000 gain

----------------------------------

Q - 6

By virtue of put option, the sale price achieved will be = max (K, S)

Gain or loss = [(Sale price - cost to produce) - option hedging cost x (1 + interest rate)] x Quantity

Cost to produce = $ 1.65 / bushel

Quantity = 20,000

interest rate = 4%

S = $ 1.70 / bushel

Using put option of strike price K = $ 1.80; option hedging cost = 0.14

the sale price achieved will be = max (K, S) = max (1.80,1.70) = $ 1.80

Gain or loss = [(Sale price - cost to produce) - option hedging cost x (1 + interest rate)] x Quantity = [(1.80 - 1.65) - 0.14 x (1 + 4%)} x 20,000 = $ 88 gain

Using put option of strike price K = $ 1.75; option hedging cost = 0.12

the sale price achieved will be = max (K, S) = max (1.75,1.70) = $ 1.75

Gain or loss = [(Sale price - cost to produce) - option hedging cost x (1 + interest rate)] x Quantity = [(1.75 - 1.65) - 0.12 x (1 + 4%)] x 20,000 = $ 496 loss

Hence, maximum profit or minimum loss = $ 88 gain.

hence, the correct answer is option B. $ 88 gain

Q - 7

Gain / Loss = gain / loss on sale in spot market + gain / loss on short call option position

= (Spot sale price - cost price) x Quantity + [C x (1 + interest rate) - max (S - K, 0)] x Quantity = (1.60 - 1.65) x 20,000 + [0.15 - max (1.60 - 1.80, 0)] x 20,000 = $ 2,000

Hence, correct answer is option B. $ 2,000